Find the quantum number that characterizes the earth's orbit around the sun. The earth's mass is 6.0 x 1024 kg, its orbital radius is 1.5 x 1011 m, and its orbital speed is 3.0 x 104 m/s.


1
Expert's answer
2021-11-24T19:12:59-0500

According to Bohr's model:


"mvr=n\frac{h}{2\pi},\\\space\\
n=\frac{2\pi mvr}{h}=2.56·10^{74}."